3 friends, a, b and c. b is 4 years older than a. c is 8 years younger than a. their three ages total 44. how old are a, b and c?
Answer by TimothyLamb(4379)

x = age of a
y = age of b
z = age of c
---
y = x + 4
z = x - 8
x + y + z = 44
---
put the system of linear equations into standard form
---
x - y = -4
x - z = 8
x + y + z = 44
---
copy and paste the above standard form linear equations in to this solver:
http://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 16
y = 20
z = 8
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---